OpenStudy (anonymous):

using order of operations (50 - 75) divided by 5squared - 7(4) find the value

OpenStudy (anonymous):

Parentheses first (50-75) = -25 So you get -25 / 52 - 7(4)

OpenStudy (anonymous):

using PEMDAS RIGHT

OpenStudy (anonymous):

No exponents so you go to multiplication. -7(4) = -28 Now you have -25 / 52 - 28

OpenStudy (anonymous):

Yes, you use PEMDAS

OpenStudy (anonymous):

look t my qeustion it has divided by 5 squared

OpenStudy (anonymous):

I thought it was 52, you should be more clear with a 5^2.

OpenStudy (anonymous):

sorry ifixed it

OpenStudy (anonymous):

5^2 = 25 So you'd have -25/25 - 7(4) Multiply -25/25 - 28 Divide -1 - 28 No addition Subtract -1-28=-29

OpenStudy (anonymous):

Thanks